Question: Assume that the weight loss for the first month of a diet program varies between 6 pounds and 12 pounds, and is spread evenly over the range of possibilities, so that there is a uniform distribution. Find the probability of the given range of pounds lost.
More than 11 pounds

Answer by stanbon(75887)

Assume that the weight loss for the first month of a diet program varies between 6 pounds and 12 pounds, and is spread evenly over the range of possibilities, so that there is a uniform distribution. Find the probability of the given range of pounds lost.
More than 11 pounds
---------------------
The shape of the distribution is a rectangle with area = 1.
------
For this uniform distribution (12-6)*height = 1
height = 1/6
-------
P(11<= x <=12) = 1*(1/6) = 1/6
===================================
Cheers,
Stan H.